# The Changes of the Northern Hadley Cell Strength in Reanalyses and Radiosonde Observations

### v1

@@ Abstract
This study examines mean meridional winds and their trends in the Northern Hadley Cell (NHC) from 1980 to 2022 using reanalysis datasets and radiosonde observations. Compared to radiosonde data, reanalyses underestimate the mean upper-tropospheric poleward flow of the NHC but accurately capture the mean equatorward flow in the lower troposphere. While climate models generally project a weakening of the NHC, our study finds no significant trend in radiosonde observations, adding to the uncertainty in future climate projections. In contrast, reanalyses indicate a strengthening, primarily due to an intensification of the upper-tropospheric poleward flow. Our examination of ERA5 analysis increments confirms that the NHC strengthening trend in ERA5 is not an artefact of data assimilation. Instead, the increments correct the first guess, which underestimates the strength of the NHC, nudging it toward a stronger circulation.
@@
The Hadley circulation (HC) is a key atmospheric system that transports heat and moisture between the tropics and subtropics. It consists of two symmetric cells that converge near the equator, characterized by rising air in the tropics, poleward flow in the upper troposphere, descending air in the subtropics, and lower-tropospheric trade winds returning toward the equator. This study investigates changes in meridional winds within the Northern Hadley cell (NHC) and its overall strength using weather balloon wind measurements (radiosondes) and reanalysis datasets, which reconstruct past weather conditions. While climate models project a weakening of the NHC, supported by studies based on sea-level pressure data, reanalyses suggest that the NHC has strengthened over recent decades. However, our study finds a neutral trend in NHC strength based on radiosonde data, which further contributes to the uncertainty and potential unreliability of future climate projections. Although previous research suggested that NHC strengthening in reanalyses results from data assimilation artefacts, this study shows that in ERA5, data assimilation corrects the weak NHC bias, but does not drive its strengthening trend.
@@ 1 Introduction
The Hadley circulation (HC) is a large-scale thermally driven atmospheric meridional overturning circulation that spans tropical and subtropical regions. It consists of two symmetric cells extending from the Inter-tropical convergence zone (ITCZ) to the subtropics and plays a pivotal role in the transport of moisture, heat and momentum between the tropics and higher latitudes (e.g. Peixoto and Oort, 1992).
@@ 2.4 Hadley Circulation Strength
The Hadley circulation strength is typically evaluated using the mass-weighted zonal-mean stream function $\psi$ in the latitude-pressure ($\varphi$-$p$) plane (Pikovnik et al., 2022). The stream function is computed by the vertical integration of the zonal-mean meridional wind $[v]$ as
@@
where $R$ is Earth’s radius, $g$ is gravity, $\varphi$ is latitude, and $p$ is pressure. The widely used HC strength metrics are based on single-point values (maxima or minima) of $\psi$ at a certain level or within the whole Hadley cell (Pikovnik et al., 2022). A stream function $\psi$ can be computed using the zonally-averaged mass conservation equation (Oort and Yienger, 1996). An accurate zonal average requires regularly spaced data in the periodic zonal domain. This condition is satisfied by reanalysis or climate model data, but not by non-uniform observations. Due to an insufficient number of radiosonde data to accurately evaluate the zonal-mean meridional wind for each latitude belt and pressure level, we compute a weighted average of the meridional wind data $v_{i}$ over the whole spherical area of the annual-mean NHC as
$[v]_{V}^{\mathrm{NHC}}(p)=\left(\sum_{i}w_{i}v_{i}\right)/\left(\sum_{i}w_{i}\right)\,.$ (2)
The weight $w_{i}$ assigned to each data point corresponds to the area of its associated Voronoi cell within the longterm-average annual-mean NHC boundaries (Figure 1c). Spherical polygon cells were generated using spherical Voronoi tessellation (Caroli et al., 2009), applied to the available radiosonde data at each pressure level and time instance. Polygon edges intersecting NHC boundaries were adjusted accordingly before computing the spherical area of each polygon. The averaging based on Voronoi cells is denoted as $[\cdot]_{V}$. For a fair comparison, the same procedure is applied to reanalysis data by first interpolating the reanalysis values to observation locations ($\mathcal{H}(v_{a})$) and then performing the Voronoi averaging.
@@
$\tilde{\psi}\right>=\left(\sum_{k=1}^{N}\tilde{\psi}_{k}\Delta p_{k}\right)/\left(\sum_{k=1}^{N}\Delta p_{k}\right)\,,$ (4)
@@ 3.1 Comparison of meridional winds in ERA5 and radiosonde observations
The representation of the Hadley circulation in ERA5 in terms of the meridional wind cross section is shown in Figure 2a, and the stream function representation is shown in Figure S3. On average, the annual-mean Hadley circulation in ERA5 extends from 32°S to 31°N, with the boundary between the Northern Hadley cell (NHC) and the Southern Hadley cell (SHC) located at 6°N (Pikovnik et al., 2022). Figure 2a captures the main features of the Hadley circulation, the upper-tropospheric poleward flow and the lower-tropospheric equatorward flow. Unlike the zonal-mean picture, Figure 2b, which presents ERA5 meridional winds interpolated to the observation space, reveals a different pattern. Specifically, the upper-tropospheric poleward flow in NHC appears to be stronger and extends further north in the zonal mean of the observation space. This discrepancy arises because radiosonde measurements predominantly capture the upper-tropospheric poleward flow over continents, with limited coverage of the upper-tropospheric equatorward flow over the Atlantic and Pacific (see Figure 1a).
@@
Figure 2: (a) Annual-mean zonal-mean meridional wind in ERA5 model space, (b) in ERA5 interpolated to observation space of the radiosondes, (c) in radiosonde data. The zonal averaging of meridional winds in (a,b,c) is performed for 36 latitude bins of 5 width. (d) Observed meridional wind averaged over the Northern Hadley cell (from 6°N to 31°N]) using the Voronoi tessellation method, compared to its ERA5 equivalent in the observation space and the mean meridional wind computed in the physical space, all for ERA5 during the 1980-2022 period.
@@
Radiosonde data suggest that the NHC extends even further north, reaching approximately 40°N (Figure 2c) and that the upper-tropospheric poleward flow is up to 0.5 m/s stronger than its ERA5 observation-space counterpart (Figure 2b). Conversely, observations indicate a slightly weaker equatorward return flow near the surface. In contrast to NHC, the SHC is poorly sampled by radiosondes. This leads to large discrepancies between the ERA5 model-space and observation-space zonal means, particularly in the upper troposphere, where even the sign of the mean meridional wind differs (compare Figures 2a and 2b).
@@
To mitigate the impact of inhomogeneous radiosonde data, Voronoi averaging of meridional winds was applied over the annual-mean NHC extent, ranging from 6°N to 31°N. Figure 2d compares the mean meridional wind in the physical space of ERA5 with the mean meridional wind of ERA5 in the observation space of the radiosondes. The results suggest that, despite the limited number of observation points, an advanced weighting method such as Voronoi tessellation can effectively represent the mean meridional wind profile in the physical space much better than other methods of averaging that we tested (i.e., simple observation-space averaging and binning averaging; Figure S4). This is particularly true for the lower-tropospheric equatorward flow, whereas the upper-tropospheric poleward flow tends to be overestimated. Despite that, the observations clearly indicate that ERA5 underestimates the upper-tropospheric poleward flow. However, note that, despite Voronoi-averaging, a perfect closure of the mass fluxes has not been achieved in the observation space.
@@ 3.2 Verification of data assimilation in ERA5 against radiosonde observations
The underestimation of the NHC upper-tropospheric poleward flow in the ERA5 first guess against observations (Figure S5a) suggests that the data assimilation increments should correct the bias by strengthening the circulation. Indeed, Figure 3a shows that the data assimilation system generates a systematic quadrupole pattern of zonal-mean meridional wind analysis increments, which intensify the northern and southern Hadley cells. This is further confirmed by the associated stream function of the analysis increment in Figure 3b, as well as by the slightly enhanced upper-tropospheric poleward flow and lower-tropospheric equatorward flow in the analysis (Figure S5a), in agreement with <>[their Figure S4]Mayer2025.
@@
Figure 3: (a) Annual-mean zonal-mean analysis increments of meridional wind computed in physical space; (b) annual-mean stream function (contours) and the stream function of the analysis increment (colors); (c) analysis increments of meridional wind in the radiosonde observation space; (d) first-guess departures of meridional wind, representing the difference between observations and the short-range forecast before data assimilation. All panels correspond to the ERA5 reanalysis and are averaged over the 1980-2022 period. In (b), contour lines represent the annual-mean zonal-mean Hadley circulation strength, with red contours indicating positive stream function values ($0.1$, $0.3$, $0.6$, $1$$\times$ 1011 kg s-1) and blue contours showing their negative counterparts ($-0.1$, $-0.3$, $-0.6$, $-1$$\times$ 1011 kg s-1). Mean increments are represented by color shading.
@@
To better understand this process, we also examined the analysis increments in the radiosonde observation space (Figure 3c) and compared them to the first-guess departures in the same space (Figure 3d). The results show that the data assimilation system corrects and partly offsets an underestimation of the HC strength present in the ERA5 first guess, which originates from the underlying forecast model. This provides further evidence that NHC is underestimated in ERA5 reanalyses.
@@
We further investigate whether the data assimilation system in ERA5 reanalysis contributes to the strengthening trend in NHC, as speculated by Chemke and Polvani (2019). First, we observe that the annual-mean zonal-mean first-guess departures of meridional winds have decreased over recent decades, reducing the magnitude of the analysis increments applied by data assimilation (Figures S6, S7c). Although the analysis increments tend to strengthen the circulation in recent decades, their magnitude has been decreasing (Figure S7c). This suggests that the analysis increments counteract the strengthening trend seen in the first guess (Figure S7b), leading to weaker strengthening of the (re)analysis (Figure S7a).
@@
Since the same forecast model is used throughout the reanalysis period, the reduction in analysis increments magnitude likely reflects improvements in the global observing system, leading to progressively more accurate analyses. In turn, this enhances the quality of the first-guess produced by the ERA5 forecast model, reducing first-guess departures and completing the feedback loop. Our research thus suggests that NHC strengthening in the ERA5 reanalysis is not a direct artefact of data assimilation. While our investigation of analysis increments has been restricted to meridional winds, note that data assimilation may also introduce artefacts in other variables, which could drive spurious trends in HC strength in the first guess.
@@ 3.3 NHC strength and its trends in reanalyses and observations
Analyzing trends in the mean meridional wind profile, we find that the reanalyses show a significant strengthening of the poleward flow - a trend not present in the observations (Figures 4b,d). However, trends in the mid-troposphere (500 hPa) and lower levels in the reanalyses largely align with observations, indicating a weakening of the trade winds. This is consistent with the reported decrease in the zonal-mean meridional gradient of mean-sea-level pressure (Chemke and Yuval, 2023). The differences between the upper-tropospheric meridional wind trends in the reanalyses and radiosonde observations raise concerns about the reliability of reanalyses to capture long-term climate trends in atmospheric circulation.
To further investigate these changes, we computed the strength of the Northern Hadley cell using stream-function metric (Equations 3 and 4), providing a broader perspective on circulation changes and trends. Figure 4e presents the standardized time series of the NHC strength, which shows a matching variability of the NHC strength in reanalyses and observations since around the 1990s, while there is a mismatch in the 1980s. Reanalyses consistently depict a significant strengthening (Figure 4e), whereas observations show a more nuanced picture, with an insignificant strengthening over the 1980-2015 period and a neutral trend over the 1980-2022 period. However, as shown in Figure 4e, much of the apparent strengthening trend can be attributed to the low NHC strength in the 1980s, which reanalyses also underestimate relative to observations. Similarly to the results reported in Pikovnik et al. (2022) and Zaplotnik et al. (2022), the NHC strengthening in ERA5 continued until around 2002, followed by a mild decline.
@@
Figure 4: (a,c) Mean meridional wind in the observation space averaged across the Northern Hadley cell (from 6°N to 31°N) using the Voronoi tessellation method. (b,d) Multidecadal trends of the mean meridional wind in (b) ERA5, JRA-55 and MERRA-2 reanalysis, and (d) NOAA 20CRv3 reanalysis compared to observations. (e) Standardized time series of the stream function $\langle\tilde{\psi}\rangle$ from 1980 to 2022. (f) Multidecadal trend of $\langle\tilde{\psi}\rangle$. (a,b) are shown for the period 1980-2022, and (c,d) are shown for the period 1980-2015. Shaded areas in (b,d) indicate the standard error of the trend, which depends on the time series length, the correlation coefficient between both time series, and the standard deviations of both time series. The black bars in (f) denote the standard errors of the trend estimate.

### v2

@@ Abstract
This study examines mean meridional winds and their trends in the Northern Hadley Cell (NHC) from 1980 to 2022 using reanalysis datasets and radiosonde observations. Compared to radiosonde data, reanalyses underestimate the mean upper-tropospheric poleward flow of the NHC but accurately capture the mean equatorward flow in the lower troposphere. While climate models generally project a weakening of the NHC, our study finds no significant trend in radiosonde observations, adding to the uncertainty in future climate projections. In contrast, reanalyses indicate a strengthening, primarily due to an intensification of the upper-tropospheric poleward flow. Our examination of ERA5 analysis increments confirms that the NHC strengthening trend in ERA5 is not an artefact of data assimilation. Instead, the increments correct the first-guess, which underestimates the strength of the NHC, nudging it toward a stronger circulation.
@@
The Hadley circulation (HC) is a key atmospheric system that transports heat and moisture between the tropics and subtropics. It consists of two cells that converge near the equator, characterized by rising air in the tropics, poleward flow in the upper troposphere, descending air in the subtropics, and lower-tropospheric trade winds returning toward the equator. This study investigates changes in meridional winds within the Northern Hadley cell (NHC) and its overall strength using weather balloon wind measurements (radiosondes) and reanalysis datasets, which reconstruct past weather conditions. While climate models project a weakening of the NHC, supported by studies based on sea-level pressure data, reanalyses suggest that the NHC has strengthened over recent decades. However, our study finds a neutral trend in NHC strength based on radiosonde data, which further contributes to the uncertainty and potential unreliability of future climate projections. Although previous research suggested that NHC strengthening in reanalyses results from data assimilation artefacts, this study shows that in ERA5, data assimilation corrects the weak NHC bias, but does not drive its strengthening trend.
@@ 1 Introduction
The Hadley circulation (HC) is a large-scale thermally driven atmospheric meridional overturning circulation that spans tropical and subtropical regions. It consists of two cells extending from the Inter-tropical convergence zone (ITCZ) to the subtropics and plays a pivotal role in the transport of moisture, heat and momentum between the tropics and higher latitudes (e.g. Peixoto and Oort, 1992).
@@ 2.4 Hadley Circulation Strength
The Hadley circulation strength is typically evaluated using the mass-weighted zonal-mean stream function $\psi$ in the latitude-pressure ($\varphi$-$p$) plane (Pikovnik et al., 2022). A stream function $\psi$ can be computed using the zonally-averaged mass conservation equation (Oort and Yienger, 1996), typically by the vertical integration of the zonal-mean meridional wind $[v]$ as
@@
where $R$ is Earth’s radius, $g$ is gravity, $\varphi$ is latitude, and $p$ is pressure. The widely used HC strength metrics are based on single-point values (maxima or minima) of $\psi$ at a certain level or within the whole Hadley cell (Pikovnik et al., 2022). Accurately computing the zonal-mean meridional wind requires data that are regularly spaced in the periodic zonal domain, as found in reanalysis or climate model outputs. To obtain the zonal average from non-uniform observational data, we compute a weighted average of the meridional wind values $v_{i}$ within each latitude belt, as follows
$[v]_{V}(\varphi,p)=\left(\sum_{i}w_{i}v_{i}\right)/\left(\sum_{i}w_{i}\right)\,.$ (2)
The weight $w_{i}$ assigned to each data point corresponds to the area of its associated Voronoi cell within a given latitude belt. Figure 1c illustrates how Voronoi cells are used to compute the mean meridional wind over the entire spherical region bounded by the long-term annual-mean NHC extent, denoted as $[v]_{V}^{\mathrm{NHC}}$. The Voronoi cells are generated using spherical Voronoi tessellation (Caroli et al., 2009) based on radiosonde data at each pressure level and time instance. Polygon edges intersecting NHC boundaries are adjusted prior to computing the spherical area of each cell. We denote Voronoi-based averaging as $[\cdot]_{V}$. For a fair comparison, the same method is applied to reanalysis data by first interpolating the reanalysis values to observation locations ($\mathcal{H}(v_{a})$), followed by Voronoi averaging.
@@
$\overline{\tilde{\psi}}=\left(\sum_{k=1}^{N}\tilde{\psi}_{k}\Delta p_{k}\right)/\left(\sum_{k=1}^{N}\Delta p_{k}\right)\,,$ (4)
@@ 3.1 Comparison of meridional winds in ERA5 and radiosonde observations
The representation of the Hadley circulation in ERA5 in terms of the meridional wind cross section is shown in Figure 2a, and the stream function representation is shown in Figure S4. On average, the annual-mean Hadley circulation in ERA5 extends from 32°S to 31°N, with the boundary between the Northern Hadley cell (NHC) and the Southern Hadley cell (SHC) located at 6°N (Pikovnik et al., 2022). Figure 2a captures the main features of the Hadley circulation, the upper-tropospheric poleward flow and the lower-tropospheric equatorward flow. Unlike the zonal-mean picture, the ERA5 meridional winds interpolated to the observation space and averaged using the Voronoi tessellation method reveal a somewhat different pattern (Figure 2b). Specifically, the upper-tropospheric poleward flow in NHC extends further north in the zonal mean of the observation space. This discrepancy arises mainly because of the limited coverage of the upper-tropospheric equatorward flow over the Atlantic (see Figure 1a). Despite the concentration of radiosonde sites over the continents, the Voronoi averaging ensures that the zonal-mean in the observation space adequately estimates the zonal-mean meridional winds and does not relate to zonal-mean meridional winds over land (Figure S5).
@@
Figure 2: (a) Annual-mean zonal-mean meridional wind in ERA5 model space, (b) in ERA5 observation space, and (c) in radiosonde data. The zonal averaging of meridional winds in (b,c) is performed using the Voronoi tessellation method. (d) Meridional winds from (a,b,c), meridionally averaged over the extent of the Northern Hadley cell (from 6°N to 31°N) during the 1980-2022 period.
@@
Radiosonde data suggest that the NHC extends even further north (Figure 2c) and that the upper-tropospheric poleward flow is up to 0.5 m/s stronger than its ERA5 observation-space counterpart (Figure 2b). Conversely, the radiosonde data show a somewhat weaker equatorward flow near the surface. In contrast to the NHC, the SHC is poorly sampled by radiosondes. This leads to large discrepancies between the zonal-mean values in the ERA5 model-space and observation-space zonal means, especially in the upper troposphere, where the magnitude of the mean meridional wind is severely underestimated (see Figures 2a and 2b). SHC is therefore not part of our study.
@@
To mitigate the effects of inhomogeneous radiosonde data, a Voronoi averaging of the meridional winds was applied over the annual-mean NHC extent, ranging from 6°N to 31°N. Figure 2d compares the mean meridional wind in the physical space of ERA5 with the mean meridional wind of ERA5 in the observation space of the radiosondes. The results suggest that, despite the limited number of observation points, an advanced weighting method such as Voronoi tessellation can effectively represent the mean meridional wind profile in the physical space much better than other averaging methods we tested (i.e., simple averaging in observation space and binning averaging; Figure S6). This is particularly true for the equatorward flow in the lower troposphere, while the poleward flow in the upper troposphere tends to be overestimated. Nevertheless, the observations clearly show that ERA5 underestimates the poleward flow in the upper troposphere. However, it should be noted that despite Voronoi averaging, perfect closure of the mass fluxes in the observation space was not achieved, particularly for the radiosonde data (see Figures S7 and S8).
@@ 3.2 Verification of data assimilation in ERA5 against radiosonde observations
The underestimation of the NHC upper-tropospheric poleward flow in the ERA5 first-guess against observations (Figure S9a) suggests that the data assimilation increments should correct the bias by strengthening the circulation. Indeed, Figure 3a shows that the data assimilation system generates a systematic quadrupole pattern of zonal-mean meridional wind analysis increments. As demonstrated by the stream function of the analysis increments, the data assimilation intensifies both the northern and southern Hadley cell (Figure 3b), and leads to enhanced upper-tropospheric poleward flow and lower-tropospheric equatorward flow in the ERA5 analysis compared to its first-guess (Figure S9a), in agreement with Mayer et al. (2025, their Figure S4).
@@
Figure 3: (a) Annual-mean zonal-mean analysis increments of meridional wind computed in physical space; (b) annual-mean stream function (contours) and the stream function of the analysis increment (colors); (c) analysis increments of meridional wind in the radiosonde observation space; (d) first-guess departures of meridional wind, representing the difference between observations and the short-range forecast before data assimilation. All panels correspond to the ERA5 reanalysis and are averaged over the 1980-2022 period. The zonal averaging in (c,d) is performed using the Voronoi tessellation method. In (b), contour lines represent the annual-mean zonal-mean Hadley circulation strength, with red contours indicating positive stream function values ($0.1$, $0.3$, $0.6$, $1$$\times$ 1011 kg s-1) and blue contours showing their negative counterparts ($-0.1$, $-0.3$, $-0.6$, $-1$$\times$ 1011 kg s-1). Mean increments are represented by color shading.
@@
To better understand this process, we also examined the analysis increments in the radiosonde observation space (Figure 3c) and compared them to the first-guess departures in the same space (Figure 3d). The results show that the data assimilation system corrects the underestimation of the HC strength present in the ERA5 first-guess, which originates from the underlying forecast model, but only partly compensates for it. This provides further evidence that NHC is underestimated in ERA5 reanalysis.
@@
We further investigate whether the data assimilation system in ERA5 reanalysis contributes to the strengthening trend in NHC, as speculated by Chemke and Polvani (2019). First, we observe that the annual-mean zonal-mean first-guess departures of meridional winds have decreased over recent decades, reducing the magnitude of the analysis increments applied by data assimilation (Figures S10, S11c). Although the analysis increments tend to strengthen the circulation in recent decades, their magnitude has been decreasing (Figure S11c). This suggests that the analysis increments counteract the strengthening trend seen in the first-guess (Figure S11b), leading to weaker strengthening of the (re)analysis (Figure S11a).
@@
Since the same forecast model is used throughout the reanalysis period, the reduction in the magnitude of analysis increments likely reflects improvements in the global observing system, leading to progressively more accurate analyses. In turn, this enhances the quality of the first-guess produced by the ERA5 forecast model, reducing first-guess departures and completing the feedback loop. Our research thus suggests that NHC strengthening in the ERA5 reanalysis is not a direct artefact of data assimilation. While our investigation of analysis increments has been restricted to meridional winds, note that data assimilation may also introduce artefacts in other variables, which could drive spurious trends in HC strength in the first-guess.
@@ 3.3 NHC strength and its trends in reanalyses and observations
Analyzing trends in the mean meridional wind profile, we find that the reanalyses show a significant strengthening of the poleward flow (Figures 4b,d). The trends in the mid-troposphere (500 hPa) and lower levels in the reanalyses also largely align with observations, indicating a weakening of the trade winds. This is consistent with the reported decrease in the zonal-mean meridional gradient of mean-sea-level pressure (Chemke and Yuval, 2023).
To further investigate these changes, we computed the strength of the Northern Hadley cell using stream-function metric (Equations 3 and 4), providing a broader perspective on circulation changes and trends. Figure 4e presents the standardized time series of the NHC strength, which shows a matching variability of the NHC strength in reanalyses and observations since around the 1990s, while there is a mismatch in the 1980s. In the observation space, the reanalyses consistently depict a significant strengthening (Figure 4e), whereas observations show a more nuanced picture, with an insignificant strengthening over the 1980-2015 period and a neutral trend over the 1980-2022 period. However, as shown in Figure 4e, much of the apparent strengthening trend can be attributed to the low NHC strength in the 1980s, which reanalyses also underestimate relative to observations. Similarly to the results reported in Pikovnik et al. (2022) and Zaplotnik et al. (2022), the NHC strengthening in ERA5 continued until around 2002, followed by a mild decline.
@@
Figure 4: (a,c) Mean meridional wind in the observation space averaged across the Northern Hadley cell (from 6°N to 31°N) using the Voronoi tessellation method. (b,d) Multidecadal trends of the mean meridional wind in (b) ERA5, JRA-55 and MERRA-2 reanalysis, and (d) NOAA 20CRv3 reanalysis compared to observations. (e) Standardized time series of the stream function $\overline{\tilde{\psi}}$ from 1980 to 2022. (f) Multidecadal trend of $\overline{\tilde{\psi}}$. (a,b) are shown for the period 1980-2022, and (c,d) are shown for the period 1980-2015. Shaded areas in (b,d) indicate the standard error of the trend, which depends on the time series length, the correlation coefficient between both time series, and the standard deviations of both time series. The black bars in (f) denote the standard errors of the trend estimate.